OpenStudy (anonymous):

What is the solution of the equation? -10 + sqrt x+8 = -4 a. 44 b. 36 c-2 d.28

OpenStudy (anonymous):

V(x + 8) = 10 - 4 = 6. Square the 2 sides (x + 8) = 36 -> x = .......

OpenStudy (anonymous):

You mean \[\sqrt{x}\] in present case we have sqrt x=-2 so X is a complex number 2i or -2i

OpenStudy (anonymous):

I'm so lost. My book doesnt teach me how to square root things but i think i can figure it out

OpenStudy (anonymous):

\[-10 + \sqrt x + 8 = -4\] Add 10 on both sides \[\sqrt x + 8 = 6\] Minus 8 from both sides \[\sqrt x = -2\] Square both sides \[x = (-2)^2\] Then \[x = 4\]